Two cyclists leave a city at the same time, one going east and the other going west. the west bound cyclist at 18mph faster than the east bound cyclist. after 3 hours they are 108 miles apart. how fast is each cyclist riding?
Answer by mananth(16946)

East bound x mph
west bound x+ 18
..
They are going away from each other.

So add up their speed.

combined speed = x+x + 18
(2x+18)
Time = 3 hours
Distance =108 miles
Distance = speed * time
(2x+18)*3=108
6x+54=108
6x=108-54
6x=54
/6
x=9 mph East bound
west bound 9+18=27 mph